OpenStudy (anonymous):

Part 2: Provide a system of TWO equations in slope-intercept form with no solutions. Using complete sentences, explain why this system has no solutions.

OpenStudy (amistre64):

y=mx + b y=mx + c

OpenStudy (amistre64):

where b and c are not the same

OpenStudy (anonymous):

tnk u

OpenStudy (anonymous):

<3

OpenStudy (amistre64):

yw :)

OpenStudy (amistre64):

y-mx = b y-mx = c since b not= c this is a contradiction